# MAIN: Mutual Alignment Is Necessary for instruction tuning

### v1

@@ Abstract
Instruction tuning has enabled large language models (LLMs) to achieve remarkable performance, but its success heavily depends on the availability of large-scale, high-quality instruction–response pairs. However, current methods for scaling up data generation often overlook a crucial aspect: the alignment between instructions and responses. We hypothesize that high-quality instruction–response pairs are not defined by the individual quality of each component, but by the extent of their alignment with each other. To address this, we propose a Mutual Alignment Framework (MAIN) that ensures coherence between the instruction and response through mutual constraints. Experiments demonstrate that models such as LLaMA and Mistral, fine-tuned within this framework, outperform traditional methods across multiple benchmarks. This approach underscores the critical role of instruction-response alignment in enabling scalable and high-quality instruction tuning for LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated unprecedented capabilities in comprehending human intent and performing cross-task generalization through contextual learning(Brown et al., 2020). A key breakthrough in aligning model behaviors with human expectations is primarily attributed to instruction tuning, a supervised learning paradigm that bridges the gap between pre-trained models’ latent knowledge and explicit task requirements (Ouyang et al., 2022). Through multi-task training on (instruction, response) pairs, this approach enables systematic knowledge elicitation while maintaining task-agnostic generalization (Chung et al., 2024). The effectiveness of this process is significantly influenced by the availability of high-quality instruction-response pairs at scale. In essence, the quality of data used in instruction tuning is critical to determining the performance and overall effectiveness of the model.
@@
The interdependence between instructions and responses introduces a dual-variable optimization problem, where optimizing one requires simultaneous consideration of the other, as neither can be fully optimized in isolation. Drawing inspiration from the alternating update strategy used in Expectation-Maximization (EM) algorithms (Moon, 1996), we propose a method to synthesize high-quality data, called MAIN. This framework iteratively optimizes both instructions and responses, enhancing their mutual alignment. Through this co-adaptive process, the alignment of instruction-response pairs improves progressively. We believe that the aligned pairs can significantly benefit model’s capabilities. Furthermore, we propose a simple but effective filtering strategy, mutual filter, which selects pairs with superior alignment, ultimately boosting the quality of the fine-tuning dataset.
To validate the effectiveness of our proposed MAIN, we test model with our generated instruction-tuning data on multiple benchmark tasks. Our experiments demonstrate significant improvements in output quality, instruction-following, and reasoning ability. Specifically, for the LLaMA-2-7B model, our generated data yield a 5.85% improvement in output quality compared to the most competitive baseline, Dog Instruct (Chen et al., 2024), and a 3.60% increase in instruction following ability over Better Alignment (Nguyen et al., 2024). Experimental results show that our method substantially boosts the model’s capabilities in various areas, surpassing traditional approaches. In summary, our main contributions are as follows:
@@
  • Propose a mutual alignment framework that reinforces the inner connection between instructions and responses, and develop an efficient data filtering method based on this framework.
  • Extensive experiments on multiple benchmark datasets show that our approach outperforms existing methods in enhancing instruction tuning effectiveness.
Figure 2: An overview of the data synthesis process, including mutual alignment, data augmentation, and data curation steps, aimed at creating high-quality, well-aligned instruction-response pairs from both seed and unlabeled data.
@@ 2.2 Data Synthesis Framework: MAIN
We present our data synthesis process as shown in the Figure 2. Our Approach assumes access to a base language model, a small set of high-quality seed data, and a large collection of unlabeled responses. The overall preocess performs three core steps:  Mutual Alignment, Data Augmentation, Data Curation.
@@
  • Data Augmentation: This step is to generate candidate pairs from unlabeled data, which leverages the trained reverse model $M_{r}:=p(I|R)$ to generate predicted instructions based on the unlabeled data $\mathcal{D}_{\text{unlabeled}}$, obtaining the augmented data pairs $\mathcal{D}_{\text{aug}}=\{R_{u},\hat{I}^{\prime}\}$.
  • Data Curation: This step involves data filtering to obtain the curated, well-aligned pairs. In detail, we leverage our proposed mutual-filter mechanism to select high-quality samples from the augmented pairs $\mathcal{D}_{\text{aug}}$. These retained samples are used to form the final training dataset with seed data $\mathcal{D}_{\text{filter}}=\left[\text{filter}(\mathcal{D}_{\text{aug}}),\mathcal{D}_{\text{seed}}\right]$ for fine-tuning the base model.
@@ 2.3 Mutual Alignment
0:  Final reverse model $M_{r}^{N}$, forward model $M_{f}^{N}$
1:  Initialize forward model $M_{f}^{0}\leftarrow M_{\text{base}}$;
2:  Initialize reverse model $M_{r}^{0}\leftarrow M_{\text{base}}$;
3:  for $k=0$ to $N$ do
4:     Use $M_{r}^{k}$ to generate $\hat{I}$ from $R$ in $\mathcal{D}_{\text{seed}}$;
5:     Construct $\mathcal{D}_{f}=\{(\hat{I},R)\}\cup\mathcal{D}_{\text{seed}}$;
6:     Update $M_{f}^{k}$ on $\mathcal{D}_{f}$ by minimizing the loss $\mathcal{L}_{f}$ (Equation (2)) to obtain $M_{f}^{k+1}$;
7:     Use $M_{f}^{k+1}$ to generate $\hat{R}$ from $I$ in $\mathcal{D}_{\text{seed}}$;
8:     Construct $\mathcal{D}_{r}=\{(\hat{R},I)\}\cup\mathcal{D}_{\text{seed}}$;
9:     Update $M_{r}^{k}$ on $\mathcal{D}_{r}$ by minimizing the loss $\mathcal{L}_{r}$ (Equation (4)) to obtain $M_{r}^{k+1}$;
10:  end for
11:  Return $M_{r}^{N}$ and $M_{f}^{N}$
@@
Similarity, the reverse model is trained to capture the alignment from instruction to response as guided by the forward model. The reverse model now is updated based on the latest forward model $M_{f}^{k+1}$ that generates synthetic responses $\hat{R}$ conditioned on the seed instructions:
@@
The objective of dynamic weighting is to maintain a balanced contribution of synthetic and seed data, thereby improving the alignment between the forward and reverse models. The weighting parameter $\alpha\in[0,1]$ is critical in this process, as it determines the relative influence of synthetic and seed data during fine-tuning. A static weighting scheme may result in suboptimal alignment: excessive reliance on synthetic data introduces noise, whereas an overemphasis on seed data can limit the model’s generalization ability. To mitigate this, we employ a dynamic adjustment mechanism that updates $\alpha$ based on the relative loss contributions of synthetic and seed data at each step. Specifically, for forward model alignment, $\alpha$ is updated as: Å
@@
This formulation adaptively adjusts the weight of synthetic data based on its relative loss, ensuring a balance between learning from synthetic examples and maintaining stability with seed data. By dynamically adjusting $\alpha$, the model effectively integrates new information from synthetic pairs while maintaining consistency with the high-quality seed data.
@@ 2.4 Data Augmentation
After optimizing the mutual alignment between instructions and responses, we expand our training data by generating synthetic instructions for unlabeled responses. Specifically, for each unlabeled response $R_{u}\in\mathcal{D}_{\text{unlabeled}}$, the reverse model produces a corresponding synthetic instruction $\hat{I}^{\prime}$. This yields candidate instruction–response pairs of the form $\{R_{u},\hat{I}^{\prime}\}$, which serve as approximations of how users might naturally formulate instructions for the given responses. However, not all candidate pairs are of high quality, so further curation is necessary.
@@ 2.5 Data Curation
To further improve data alignment, we introduce a simple yet effective filtering mechanism. We assume that high-quality instruction-response pairs should be well-aligned, where the predicted instruction generated by the reverse model can be decoded by the forward model to recover the response, which should closely resemble the original. This process is akin to the interaction between an encoder and a decoder (Cho et al., 2014). Thus, we select the most well-aligned pairs. Using the candidate pairs $\{R_{u},\hat{I}^{\prime}\}$ from the Augmentation stage, we then employ the forward model to generate synthetic responses $\hat{R}^{\prime}$ based on $\hat{I}^{\prime}$:
@@ 3.1 Experimental Setup
In the MAIN, we used the LLaMA-2-7B model (Touvron et al., 2023) as the base model for mutual alignment experiments, and additionally validated the generalization of our approach using the Mistral-7B-v1 model (Jiang et al., 2023). For each iteration, both the forward and reverse models are trained for one epoch. The learning rate is set to $1\times 10^{-5}$ with a linear decay schedule. The adaptation parameter, denoted as $a$, dynamically adjusts based on the formula in 5. The batch size is set to 32. For data curation, we selected the top 16,800 pairs from the unlabeled data based on cross-entropy and combined them with the seed data to create the final fine-tuning dataset.
@@
We use the pre-trained LLaMA-2-7B model and Mistral-7B-v1 model as the base models for fine-tuning respectively. Detailed hyperparameter configurations are provided in Appendix A.
3.2 Evaluation
@@
Instruction-following ability is evaluated with IFEval (Zhou et al., 2023), which reports accuracy across four metrics: Prompt-level Strict (P-S), Instruction-level Strict (I-S), Prompt-level Loose (P-L (Zhou et al., 2023), ensuring a comprehensive assessment of instruction adherence.
@@ 3.3 Baselines
We compare our framework to several baseline approaches, all evaluated on the Falcon-RefinedWeb dataset with 20K samples.
@@
This method (Li et al., 2023a) is a two-stage curation process that filters and selects high-quality instruction–response pairs before fine-tuning.
@@
This method (Nguyen et al., 2024) first generates instructions via back-translation, then filters low-quality pairs to obtain high-quality response.
Base Model Method Output quality Instruction following Reasoning ability
@@
Llama-2-7B Humpback (Li et al., 2023a) 41.02 15.46 18.39 26.42 29.91 55.90 44.91 79.42 73.32 44.48
Longform (Köksal et al., 2023) 35.64 15.23 17.56 26.10 29.29 55.72 45.02 78.98 73.21 45.07
Dog Instruct(Chen et al., 2024) 52.35 15.52 19.40 28.17 32.01 56.06 45.62 79.89 74.13 45.77
Better Alignment (Nguyen et al., 2024) 50.37 16.82 19.69 27.70 31.52 55.92 45.84 80.33 74.12 45.30
MAIN 58.20 20.22 23.36 31.17 35.37 57.08 45.47 81.22 74.51 47.40
$\Delta$ over Best Result +5.85 +3.40 +3.67 +3.00 +3.36 +1.02 -0.37 +0.89 +0.38 +1.63
@@
Table 1: Benchmarking results of different instruction tuning methods on Llama-2-7B and Mistral-7B using Falcon RefinedWeb dataset given same data quantity. $\Delta$ over Best Result quantify improvements relative to the strongest baseline method across evaluation categories.
@@ 4 Experimental Results
This section presents the experimental results, including quantitative evaluations, an ablation study, and a case study, to assess the effectiveness of our approach.
@@ 4.1 Quantitative Results
We conduct experiments across three benchmarks, each assessing a different aspect of the MAIN: AlpacaEval evaluates output quality, IFEval measures instruction-following ability, and OpenLLM tests reasoning capability.
Output Quality.
As shown in Table 1, our method achieves the highest win rate in AlpacaEval dataset, surpassing the leading baseline. Specifically, on Llama-2-7B, our method achieves a win rate of 58.20%, representing a 5.85% improvement over the best baseline Dog Instruct (Chen et al., 2024). On Mistral-7B, our method outperforms Better alignment (Nguyen et al., 2024) by 3.15%, reaching a win rate of 48.94% compared to 45.79%. These results confirm that our MAIN method enhances instruction-response alignment more effectively than previous approaches, leading to outputs that better align with human expectations.
@@
Table 1 illustrates the performance of our method in IFEval dataset, where it outperforms previous approaches. Compared to the best-performing baseline Better alignment (Nguyen et al., 2024), our approach achieves consistent improvements across all evaluation metrics. Specifically, on Llama-2-7B, we see an increase of 2.59% in P-S and 3.42% in I-S. For Mistral-7B, we observe a 5.12% improvement in P-S and a 4.84% improvement in I-S over Better alignment (Nguyen et al., 2024). These results highlight the crucial role of enhanced data alignment in fine-tuning, which allows our model to better interpret and respond to user instructions, thereby driving its superior performance in instruction-following tasks.
@@
As shown in Table 1, our method demonstrates strong improvements in reasoning and factual accuracy across multiple downstream tasks. On Llama-2-7B, our approach shows a 2.02% improvement over the best baseline, Better Alignment, on ARC-Challenge and a 1.63% improvement on TruthfulQA. In particular, ARC-Challenge benefits from our method’s ability to better capture common-sense reasoning patterns, which likely leads to more accurate responses. On Mistral-7B, the most significant improvements are observed in TruthfulQA, where our method outperforms Better Alignment by 5.03%, and in MMLU, with a 2.65% increase. These benchmarks, which require accurate factual recall and complex reasoning, show how our method strengthens the model’s ability to provide correct and contextually appropriate answers.
These results underline the effectiveness of our MAIN in enhancing both reasoning and factual accuracy. By refining the alignment of instruction and response pairs during fine-tuning, our model is better equipped to handle complex reasoning tasks and provide more precise, reliable outputs across various challenging benchmarks.

### v2

@@ Abstract
Instruction tuning has empowered large language models (LLMs) to achieve remarkable performance, yet its success heavily depends on the availability of large-scale, high-quality instruction-response pairs. To meet this demand, various methods have been developed to synthesize data at scale. However, current methods for scaling up data generation often overlook a crucial aspect: the alignment between instructions and responses. We hypothesize that the quality of instruction-response pairs is determined not by the individual quality of each component, but by the degree of mutual alignment. To address this, we propose a Mutual Alignment Framework (MAIN) which enforces coherence between instructions and responses through mutual constraints. We demonstrate that MAIN generalizes well across model architectures and sizes, achieving state-of-the-art performance on LLaMA, Mistral, and Qwen models across diverse benchmarks. This work underscores the critical role of instruction-response alignment in enabling generalizable and high-quality instruction tuning for LLMs.
@@ 1 Introduction
Large Language Models have demonstrated unprecedented capabilities in comprehending human intent and performing cross-task generalization through contextual learning(Brown et al., 2020). A key breakthrough in aligning model behaviors with human expectations is primarily attributed to instruction tuning, a supervised learning paradigm that bridges the gap between pre-trained models’ latent knowledge and explicit task requirements (Ouyang et al., 2022). Through multi-task training on (instruction, response) pairs, this approach enables systematic knowledge elicitation while maintaining task-agnostic generalization (Chung et al., 2024). The effectiveness of this process is significantly influenced by the availability of high-quality instruction-response pairs at scale. In essence, the quality of data used in instruction tuning is critical to determining the performance and overall effectiveness of the model.
@@
The interdependence between instructions and responses introduces a dual-variable optimization problem, where enhancing one component necessitates adjusting the other simultaneously, as neither can be optimized in isolation. Drawing inspiration from the alternating update strategy used in Expectation-Maximization (EM) algorithms (Moon, 1996), we propose MAIN, a framework for synthesizing high-quality data. This framework iteratively optimizes both instructions and responses, progressively reinforcing their mutual alignment. Through this co-adaptive process, the alignment between instruction-response pairs improves substantially, which we believe will significantly boost the model’s performance. Furthermore, we propose a straightforward but effective filtering strategy, mutual filter, which selects pairs with superior alignment, ultimately boosting the quality of the fine-tuning dataset.
To validate the effectiveness of our proposed MAIN framework, we conducted extensive evaluations by fine-tuning models with instruction-response pairs generated by MAIN across multiple benchmarks. Experimental results demonstrate substantial improvements in output preference, instruction-following capability, and reasoning ability. Specifically, for the LLaMA-2-7B model, our framework achieves a 5.85% increase in output preference compared to Dog Instruct (Chen et al., 2024), and a 3.60% improvement in instruction-following ability over Better Alignment (Nguyen et al., 2024). Furthermore, additional analyses, including experiments on filtering strategies and GPT-4-based pairwise evaluations of instruction alignment, confirm that MAIN’s mutual alignment enhances the coherence and quality of instruction-response pairs. Our primary contributions are as follows:
@@
  • We propose MAIN, a mutual alignment framework that reinforces the inner connection between instructions and responses, and develop a straightforward but efficient data filtering method.
  • We conduct extensive evaluations across diverse model families and parameter scales, showing that MAIN outperforms existing methods in enhancing instruction tuning effectiveness.
@@ 2 Methodology
Figure 2: An overview of the data synthesis process, including mutual alignment, data augmentation, and data curation, aimed at creating high-quality, well-aligned instruction-response pairs from both seed and unlabeled data.
@@ 2.2 Data Synthesis Framework: MAIN
We present our data synthesis framework, MAIN, illustrated in Figure 2. Given a base language model, a small set of high-quality seed pairs, and a large collection of unlabeled responses, MAIN constructs a high-quality training dataset through three tightly coupled stages: Mutual Alignment, Data Augmentation, and Data Curation.
@@
  • Data Augmentation: With the reverse model $M_{r}$ trained in the previous step, we apply it to the unlabeled response data $\mathcal{D}_{\text{unlabeled}}$ to generate corresponding pseudo-instructions. This yields a set of candidate pairs $\mathcal{D}_{\text{aug}}=\{(R_{u},\hat{I}^{\prime})\}$, expanding the data space beyond the original seed data.
  • Data Curation: Not all augmented pairs are equally reliable. To select high-quality examples, we apply our mutual filter, which uses both the forward and reverse models to assess alignment consistency. Only examples that meet mutual alignment criteria are retained. These filtered pairs, combined with the original seed data, form the final fine-tuning dataset: $\mathcal{D}_{\text{filter}}=\text{filter}(\mathcal{D}_{\text{aug}})\cup\mathcal{D}_{\text{seed}}$.
@@ 2.3 Mutual Alignment
0:  Reverse model $M_{r}^{N}$, forward model $M_{f}^{N}$
1:  Initialize $M_{f}^{0}\leftarrow M_{\text{base}}$, $M_{r}^{0}\leftarrow M_{\text{base}}$
2:  for $k=0$ to $N-1$ do
3:     Generate $\hat{I}$ from $R$ in $\mathcal{D}_{\text{seed}}$ using $M_{r}^{k}$
4:     Build training set $\mathcal{D}_{f}=\{(\hat{I},R)\}\cup\mathcal{D}_{\text{seed}}$
5:     Update $M_{f}^{k}$ on $\mathcal{D}_{f}$ by minimizing loss $\mathcal{L}_{f}$ (Equation (2)) to obtain $M_{f}^{k+1}$
6:     Generate pseudo-responses $\hat{R}$ from $I$ in $\mathcal{D}_{\text{seed}}$ using $M_{f}^{k+1}$
7:     Build training set $\mathcal{D}_{r}=\{(\hat{R},I)\}\cup\mathcal{D}_{\text{seed}}$
8:     Update $M_{r}^{k}$ on $\mathcal{D}_{r}$ by minimizing loss $\mathcal{L}_{r}$ (Equation (4)) to obtain $M_{r}^{k+1}$
9:  end for
10:  Return $M_{r}^{N}$ and $M_{f}^{N}$
@@
Similarly, the reverse model is trained to capture the alignment from instruction to response as guided by the forward model. The reverse model now is updated based on the latest forward model $M_{f}^{k+1}$ that generates synthetic responses $\hat{R}$ conditioned on the seed instructions:
@@
To balance the influence of synthetic and seed data, we adopt a dynamic weighting strategy that adaptively adjusts their contributions during training. The weighting coefficient $\alpha\in[0,1]$ controls this balance, where static settings may lead to suboptimal outcomes: over-reliance on synthetic data can introduce noise, while overemphasis on seed data may hinder generalization. To address this, we update $\alpha$ at each step based on the relative loss of the two data sources. For forward model training, the update rule is:
@@
This formulation ensures that the contribution of synthetic data is modulated according to its training loss, enabling the model to incorporate novel patterns from synthetic pairs while preserving the stability offered by seed data.
@@ 2.4 Data Augmentation
Following mutual alignment optimization, we expand the training corpus by generating synthetic instructions for unlabeled responses. Specifically, for each unlabeled response $R_{u}\in\mathcal{D}_{\text{unlabeled}}$, the reverse model generates a corresponding synthetic instruction $\hat{I}^{\prime}$, forming a set of candidate instruction–response pairs $\{(R_{u},\hat{I}^{\prime})\}$. These pairs approximate how users might naturally formulate instructions for the given responses. However, as the quality of generated pairs may vary, a subsequent curation step is required to ensure alignment consistency and data reliability.
@@ 2.5 Data Curation
To further improve data alignment, we introduce an effective filtering mechanism. We assume that high-quality instruction-response pairs should be well-aligned, where the predicted instruction generated by the reverse model can be decoded by the forward model to recover the response, which should closely resemble the original. This process is akin to the interaction between an encoder and a decoder (Cho et al., 2014). Thus, we select the most well-aligned pairs. Using the candidate pairs $\{R_{u},\hat{I}^{\prime}\}$ from the Augmentation stage, we then employ the forward model to generate synthetic responses $\hat{R}^{\prime}$ based on $\hat{I}^{\prime}$:
@@ 3.1 Experimental Setup
For mutual alignment experiments, we adopt LLaMA-2-7B as the base model, and additionally evaluate the generalization of our approach on Mistral and Qwen models. In each iteration, both the forward and reverse models are trained for one epoch. We use a learning rate of $1\times 10^{-5}$ with a linear decay schedule. The adaptation weight $\alpha$ is dynamically updated according to Equation 5. The batch size is set to 32. For data curation, we select the top 16,800 instruction-response pairs from the unlabeled set based on cross-entropy scores and combine them with the seed data to form the final fine-tuning dataset.
@@
We use the pre-trained LLaMA, Mistral and Qwen as the base models respectively. Detailed hyperparameter configurations are provided in Appendix A.
3.2 Benchmarks
@@
Instruction-following ability is evaluated with IFEval (Zhou et al., 2023), which reports accuracy across four metrics: Prompt-level Strict (P-S), Instruction-level Strict (I-S), Prompt-level Loose (P-L), Instruction-level Loose (I-L) ensuring a comprehensive assessment of instruction adherence.
@@ 3.3 Baselines
We compare our framework to several baseline approaches, fine-tuned on 3.2k seed data and 16.8k generated data from Falcon-RefinedWeb dataset.
@@
This method (Li et al., 2023a) is a two-stage curation process that filters and selects high-quality pairs before fine-tuning.
@@
This method (Nguyen et al., 2024) generates instructions via back-translation, then filters pairs to obtain high-quality response.
Base Model Method Output preference Instruction following Reasoning ability
@@
Llama-2-7B Humpback (Li et al., 2023a) 41.02 15.46 26.42 18.39 29.91 55.90 44.91 79.42 73.32 44.48
Longform (Köksal et al., 2023) 35.64 15.23 26.10 17.56 29.29 55.72 45.02 78.98 73.21 45.07
Dog Instruct(Chen et al., 2024) 52.35 15.52 28.17 19.40 32.01 56.06 45.62 79.89 74.13 45.77
Better Alignment (Nguyen et al., 2024) 50.37 16.82 27.70 19.69 31.52 55.92 45.84 80.33 74.12 45.30
MAIN 58.20 20.22 31.17 23.36 35.37 57.08 45.47 81.22 74.51 47.40
$\Delta$ over Best Result +5.85 +3.40 +3.00 +3.67 +3.36 +1.02 -0.37 +0.89 +0.38 +1.63
@@
Qwen2.5-14B Humpback (Li et al., 2023a) 49.83 69.45 79.92 74.28 82.25 65.25 77.04 83.81 78.91 57.63
Longform (Köksal et al., 2023) 45.42 69.12 79.37 74.05 82.46 65.12 76.92 83.04 76.55 55.72
Dog Instruct (Chen et al., 2024) 50.27 72.83 81.89 76.64 84.72 64.71 78.33 84.19 79.12 56.34
Better Alignment (Nguyen et al., 2024) 53.93 71.37 81.55 76.32 84.14 65.36 78.01 84.25 79.63 55.91
MAIN 61.30 74.90 84.03 79.93 87.61 66.93 80.26 84.07 80.72 59.49
$\Delta$ over Best Result +7.37 +2.07 +2.14 +2.29 +2.89 +1.57 +1.93 -0.18 +1.09 +3.15
Table 1: Benchmarking results of different methods on Llama-2-7B, Mistral-7B and Qwen2.5-14B using Falcon RefinedWeb dataset given same data quantity (20k samples). $\Delta$ over Best Result quantify improvements relative to the strongest baseline method across evaluation categories.
@@ 4 Experimental Results
This section presents the experimental results, including quantitative results, an ablation study, data alignment analysis and a case study, to assess the effectiveness of our approach.
@@ 4.1 Quantitative Results
We conduct experiments across three benchmarks, each assessing a different aspect of the MAIN: AlpacaEval evaluates output preference, IFEval measures instruction-following ability, and OpenLLM tests reasoning capability, with additional experiments detailed in Appendix B
Output Preference.
As shown in Table 1, our method achieves the highest win rate in AlpacaEval dataset, surpassing the leading baseline. Specifically, on Llama-2-7B, our method achieves a win rate of 58.20%, representing a 5.85% improvement over the best baseline Dog Instruct (Chen et al., 2024). On Mistral-7B, our method outperforms Better alignment (Nguyen et al., 2024) by 3.15%, reaching a win rate of 48.94% compared to 45.79%. On Qwen2.5-14B, MAIN achieves the highest win rate of 71.30%. These results confirm that our MAIN method enhances instruction-response alignment more effectively than previous approaches, leading to outputs that better align with human expectations.
@@
Table 1 illustrates the results of our method in IFEval dataset where MAIN achieves state-of-the-art performance across all three model backbones. Compared to the best-performing baseline, our approach achieves consistent improvements across all evaluation metrics. Specifically, on Llama-2-7B, we see an increase of 2.59% in P-S and 3.42% in I-S. For Mistral-7B, we observe a 5.12% improvement in P-S and a 4.84% improvement in I-S over Better alignment (Nguyen et al., 2024). For Qwen2.5-14B, MAIN also leads with a +2.07% and +2.99% gain in P-S and I-S respectively. These results highlight the crucial role of enhanced data alignment in fine-tuning, which allows our model to better interpret and respond to user instructions, thereby driving its superior performance in instruction-following tasks.
@@
As shown in Table 1, our method demonstrates strong improvements in reasoning and factual accuracy across multiple downstream tasks. On Llama-2-7B, our approach shows a 2.02% improvement over the best baseline, Better Alignment, on ARC-Challenge and a 1.63% improvement on TruthfulQA. In particular, ARC-Challenge benefits from our method’s ability to better capture common-sense reasoning patterns, which likely leads to more accurate responses. On Mistral-7B, the most significant improvements are observed in TruthfulQA, where our method outperforms Better Alignment by 5.03%, and in MMLU, with a 2.65% increase. We further include evaluation on Qwen2.5-14B, which follows a similar trend. These benchmarks that require accurate factual recall and complex reasoning, show how our method strengthens the model’s ability to provide correct and contextually appropriate answers.